OpenStudy (anonymous):

can somebody help me with this question? 2: let f be the function given by : f(x)=((x^4)-(16x^2))^(1/2) a) Find the domain of f. b) Find f '(x) c) Find the slope of the line normal to the graph of f at x = 5

OpenStudy (anonymous):

\[f(x)=\sqrt{x^4-16x^2}\] \[f'(x)=\frac{4x^3-32x}{2\sqrt{x^4-16x^2}}\] then some algebra

OpenStudy (anonymous):

domain is \[x^4-16x^2\geq 0\] \[x^2(x^2-16)\geq 0\] \[(-\infty, -4)\cup\{0\}\cup (4,\infty)\]

OpenStudy (anonymous):

algebra for derivative gives \[\frac{4x^3-32x}{2x\sqrt{x^2-16}}\] \[\frac{2x^2-16}{\sqrt{x^2-16}}\]

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

um.. excuse me but im still a lil confused a bit bout part c can u explain it to me again